OpenStudy (yaya090600):

Using the completing-the-square method, rewrite f(x) = x2 − 6x + 2 in vertex form.

OpenStudy (yaya090600):

A.) f(x) = (x − 3)^2 B.) f(x) = (x − 3)^2 + 2 C.) f(x) = (x − 3)^2 − 7 D.) f(x) = (x − 3)^2 + 9